# Population dynamics in the global coral–symbiont network under temperature variations

### v1

@@ 1 Introduction
The bleaching event causes several ecological impacts such as a decrease in its growth rate. Furthermore, bleaching events are recurrent and have become more frequent in the last decade [4]. However, warming events do not equally affect all organisms as they have different thermal tolerances which represents their capacity to grow under water temperature variation [8, 9]. Finally, the capacity to grow also depends on the symbiotic relationships of the organism. Intuitively a coral with multiple and diverse algae has a higher tolerance to warming events and thus more robust growth.
@@ 3 Data source and network
Data from the GeoSymbio [2] and a complementary database [8] were used to construct the bipartite coral-symbiont network. Geosymbio database provided information about the organisms, such as Symbiodinium type based on ITS2 sequence type, scientific name (genus and species) of host and the location (i.e., ocean region) from which the Symbiodinium specimen was collected. There is a total of 53 ocean regions in the GeoSymbio. The complementary database was used to obtain data on the thermal tolerances of the Symbiodinium type and the host. Unfortunately, the database is not complete and some organism do not have a specified thermal tolerance. In such cases, the mean value of the thermal tolerance was used as reference.
@@
Moreover, five connected components of the global network were chosen to be analyzed separately (Fig. 1) which represent the most threatened regions of coral bleaching in the oceans.
@@
Region Symbiont nodes Host nodes Edges
Great Barrier Reef 76 198 415
Phuket 36 152 442
Western Indian 43 131 337
Western Caribbean 36 61 111
Florida 26 32 75
@@ 3.1 Degree distribution
The degree of the global coral-symbiont network is analyzed, considering all connected components. Table 2 shows that the average degree of both types of nodes is similar. Furthermore, since the minimum degree is 1, there are no isolated organisms (nodes) in the coral-symbiont network.
@@
Fig. 2 shows the complementary cumulative distribution function (CCDF) for the degree of both symbiont and host nodes. Note that both distributions are heavy-tailed since a tiny number of symbionts are connected to more than 100 hosts and a tiny number of hosts are connected to more than 50 symbionts. Further, note that the symbionts have a heavier tail (the distribution curve decreases more slowly) indicating that symbionts connect more, also because the much larger number of hosts.
@@ 4 Mathematical model for population growth
The division of the population of an organism by its degree assumes that each population interacts uniformly with the population of neighboring organism. This assumption significantly simplifies the model.
@@
$m_{i}^{s}=\begin{cases}\mu,&\text{ if $T(t)\leq z$}\\ \displaystyle 1-e^{\left(\frac{-(T(t)-z)^{2}}{\left(\tau_{i}^{s}\right)^{2}}\right)},&\text{ if $T(t)>z$}\end{cases}$ (3)
@@ 5.1 Population growth
Fig. 4 shows the population growth for symbionts and hosts at the Great Barrier Reef region. Note that all species showed an overall increasing trend in the population. Moreover, when water temperature is far from the optimal the population of most species decreases. This same trend was observed in all other regions. However, all species showed resilience as they continue to grow in population despite the thermal stress events.
@@
Nevertheless, even with the initial populations being the same for all species, differences in the evolution of populations occur due to the structure of the network and thermal resistances. Since the network structure is not uniform, as the node degrees are very different, population growth is also not uniform. Fig. 5 shows the population distribution after 1500 days for both symbionts and hosts for all regions analyzed. Note that population distribution for hosts exhibits a heavy tail in all regions. This highlights the central role that the network structure plays in the the population dynamics.
@@ 5.2 Influence of network structure
In order to study the role of the network structure in the growth process, a random bipartite network was created for each region (Tab. 1) using a previously described methodology [8]. In particular, each original edge was repositioned uniformly at random, destroying any biological symbiotic affinity. Note that the number of edges of each network was preserved. Fig. 6 shows the population distribution for each region when growing on the random networks. Interestingly, the distributions have a much lighter tail in comparison to the original networks (see Fig. 5).
The role of the network structure and the the thermal tolerances of species on the population can be studied through correlation analysis. Table 5 shows the correlation between the thermal tolerances of species and their final populations. Note that this value is relatively low across all regions (close to zero). Therefore, final population are not even moderately correlated with the thermal tolerance.
@@
Table 6: Correlation between final populations and node degrees. For the random network, the sample average correlation and its standard deviation is reported using 20 independent instances of the random network.
@@
Great Barrier Reef     0.384        0.552   0.208 $\pm$ 0.101   0.209 $\pm$ 0.075
Phuket     0.193        0.186   0.151 $\pm$ 0.101   0.075 $\pm$ 0.080
Western Indian     0.183        0.206   0.170 $\pm$ 0.127   0.134 $\pm$ 0.081
Western Caribbean     0.469        0.579   0.176 $\pm$ 0.171   0.141 $\pm$ 0.137
Florida     0.383        0.404   0.126 $\pm$ 0.206   0.001 $\pm$ 0.165
@@
Table 6 also shows the correlation between degree and final population in the random networks. The average correlation is considerably smaller than the values observed in the original networks with moderate correlation (above 0.3), since randomly repositioning the edges removes the heavy tail property and makes the degree distribution more centered. These results reinforce the importance of network structure for the survival of these organisms.
Table 7: Correlation between final populations and degree of neighbors.

### v2

@@ 1 Introduction
The bleaching event causes several ecological impacts such as a decrease in its growth rate. Furthermore, bleaching events are recurrent and have become more frequent in the last decade [4]. However, warming events do not equally affect all organisms, as they have different thermal tolerances which represents their capacity to grow under water temperature variation [8, 9]. Finally, the capacity to grow also depends on the symbiotic relationships of the organism. Intuitively a coral with multiple and diverse algae has a higher tolerance to warming events and thus more robust growth.
@@ 3 Data source and network
Data from the GeoSymbio [2] and a complementary database [8] were used to construct the bipartite coral-symbiont network. Geosymbio database provided information about the organisms, such as Symbiodinium type based on ITS2 sequence type, scientific name (genus and species) of coral and the location (i.e., ocean region) from which the Symbiodinium specimen was collected. There is a total of 53 ocean regions in the GeoSymbio. The complementary database was used to obtain data on the thermal tolerances of the Symbiodinium type and the coral host. Unfortunately, the database is not complete and some organism do not have a specified thermal tolerance. In such cases, the mean value of the thermal tolerance was used as reference.
@@
Moreover, five connected components of the global network each corresponding to a different region were chosen to be analyzed separately, as shown in Table 1. These regions represent the most threatened regions of coral bleaching in the oceans. Note that these networks have different number of nodes and edges, but relatively similar edge density.
@@
Region Symbiont nodes Host nodes Edges Density
Great Barrier Reef 76 198 415 0.055
Phuket 36 152 442 0.162
Western Indian 43 131 337 0.120
Western Caribbean 36 61 111 0.101
Florida 26 32 75 0.180
@@ 3.1 Degree distribution
The degree of the global coral-symbiont network is analyzed, considering all 53 ocean regions. Table 2 shows that the average degree of both types of nodes is relatively similar but not the standard deviation which is larger for the symbiont nodes. Furthermore, since the minimum degree is 1, there are no isolated organisms (nodes) in the coral-symbiont network.
@@
Fig. 2 shows the complementary cumulative distribution function (CCDF) for the degree of both symbiont and host nodes. Note that both distributions are heavy-tailed since a tiny number of symbionts are connected 100 or more hosts and a tiny number of hosts are connected to 50 or more symbionts. Further, note that the symbionts have a heavier tail (the distribution curve decreases more slowly) indicating that symbionts connect more, also because the number of host nodes is much larger.
@@ 4 Mathematical model for population growth
Dividing the population of an organism by its degree assumes that each population interacts uniformly with the population of neighboring organism. This normalization ensures that the interaction of a symbiont or host population is distributed evenly among its connections. While hosts typically have fewer neighbors than symbionts, this asymmetry is inherent to the network structure and is represented in the model by this normalization. Moreover, this assumption significantly simplifies the model as it requires a single variable (population) for each node while also capturing network heterogeneity (different degrees).
@@
$m_{i}^{s}=\begin{cases}\mu,&\text{ if $T(t)\leq z$}\\ \vspace{2mm}\\ \displaystyle 1-e^{\left(\frac{-(T(t)-z)^{2}}{\left(\tau_{i}^{s}\right)^{2}}\right)},&\text{ if $T(t)>z$}\end{cases}$ (3)
@@ 5.1 Population growth
Fig. 4 shows the population growth for symbionts and hosts at the Great Barrier Reef region. Note that all species showed an overall increasing trend in the population.
@@
Moreover, when water temperature is far from the optimal the population of most species decreases. This same trend was observed in all other regions. However, all species showed resilience as they continue to grow in population despite the thermal stress events.
@@
Nevertheless, even with the initial populations being the same for all species, differences in the evolution of populations occur due to the structure of the network and thermal resistances. Since the network structure is not uniform, as the node degrees are very different, population growth is also not uniform. Fig. 5 shows the population distribution after 1500 days for both symbionts and hosts for all regions analyzed. Note that population distribution for hosts exhibits a heavy tail in all regions.This highlights the central role of network structure in population dynamics, as it determines how the species interactions shape resilience. Species with higher node degree (generalist species) adapt better to environmental changes, while species with lower node degree (specialist species) grow slower when exposed to thermal disturbances.
@@ 5.2 Influence of network structure
In order to study whether the structure of interactions (network) influences the resistance of species to thermal stresses, a random bipartite network was created for each region (as shown in Tab. 1) using a previously described methodology [8]. In particular, each original edge was repositioned uniformly at random, destroying any biological symbiotic affinity. Note that the number of edges of each network was preserved. Moreover, the network randomization procedure adopted did not allow any isolated nodes, as all nodes in the randomized network have degree of at least 1. Fig. 6 shows the population distribution for each region when growing on the random networks. Interestingly, the distributions have a much lighter tail in comparison to the original networks (see Fig. 5).
The role of the network structure and the thermal tolerances of species on the population can be studied through correlation analysis.
Table 5 shows the correlation between the thermal tolerances of species and their final populations. Note that this value is relatively low across all regions (close to zero). Therefore, final population are not even moderately correlated with the thermal tolerance.
@@
Table 6: Correlation between final populations and node degrees. For the random network, the sample average correlation and its standard deviation is reported using 50 independent instances of the random network.
@@
Great Barrier Reef     0.384        0.552   0.213 $\pm$ 0.072   0.168 $\pm$ 0.073
Phuket     0.193        0.186   0.232 $\pm$ 0.144   0.078 $\pm$ 0.087
Western Indian     0.183        0.206    0.075 $\pm$ 0.164   0.100 $\pm$ 0.092
Western Caribbean     0.469        0.579   0.154 $\pm$ 0.125   0.146 $\pm$ 0.134
Florida     0.383        0.404   0.078 $\pm$ 0.216   0.084 $\pm$ 0.202
@@
Table 6 also shows the correlation between degree and final population in the random networks. Note that the average correlation for the random networks is considerably smaller than the original networks that have moderate correlation (above 0.3). As expected, randomly repositioning the edges removes the heavy tail property and makes the degree distribution more centered. These results reinforce the importance of network structure for the survival of these organisms.
Table 7: Correlation between final populations and sum of degrees of neighbors.